OpenStudy (anonymous):

2(x+3)-1=7x with steps please

OpenStudy (anonymous):

2(x+3)-1=7x distribute the 2 into the parenthesis 2x+6-1=7x combine the like terms 2x+5=7x subtract over the 2x 5=5x Divide to get the x by itself x=1